OpenStudy (anonymous):

what is the domain of 5x^2-1

OpenStudy (anonymous):

answer choices? and heyyy welcome to open study mathmatics!!

OpenStudy (zzr0ck3r):

is there any reason that you cant plug in any real number for x?

OpenStudy (anonymous):

(f/g)(x), x^2+8x/5x^2-1, What is the domain?

OpenStudy (anonymous):

Thanks gabgurl!

OpenStudy (zzr0ck3r):

ahh much different

OpenStudy (zzr0ck3r):

ahh well the denominator cant be equal to 0 so find out what makes that true and exclude those numbers so 5x^2-1=0 implies 5x^2=1implies x^2 = 1/5 implies \(x=\pm\sqrt{\frac{1}{5}}\)

OpenStudy (zzr0ck3r):

so the domain is \(D = \{x\in \mathbb{R} \ | \ x\ne \pm\sqrt{\frac{1}{5}}\}\)

OpenStudy (anonymous):

It has to be written in interval notation.

OpenStudy (anonymous):

You are AWESOME by the way! I was close but no cigar!